Lauren and david, two cyclists leaev from philadelphia traveling in opposite directions. David travels 5mph faster than lauren, after 3 hours, they are 25miles apart, what are the rates of each?
Answer by stanbon(75887)

Lauren and david, two cyclists leave from philadelphia traveling in opposite directions. David travels 5mph faster than lauren, after 3 hours, they are 25 miles apart, what are the rates of each?
----------------------
Lauren rate = x mph
David rate = x+5 mph
-------
Combined rate = 2x+5 mph
-------
Solve rate * time = distance
(2x+5)3 = 25
6x+15 = 25
6x = 10
x = 5/3 mph
----
Note: That's awfully slow for a bicycle. Please
check the statement of your problem.
================
Cheers,
Stan H.
================